at t=0 seconds, Vikki stood on the roof of a building and threw a penny in the air. The height in feet h(t) of the ball at t seconds is given by the formula H(t) = -16t^2 + 96t + 56. a) how tall is the building? b) when is the ball at its highest point? c) to what height does the ball travel? d) when does the ball strike the ground?
Answer by stanbon(75887)

at t=0 seconds, Vikki stood on the roof of a building and threw a penny in the air. The height in feet h(t) of the ball at t seconds is given by the formula
H(t) = -16t^2 + 96t + 56.
a) how tall is the building?
H(0) = 56 feet
---------------------------
b) when is the ball at its highest point?
Maximum height occurs when t = -b/2a = -96/(2*-16) = 3 seconds
-------------------------------------------------------------------------
c) to what height does the ball travel?
Find H(3) for this answer.
-----------------------------------------
d) when does the ball strike the ground?
Solve H(t) = 0 and use the positive solution.
I get t = 6.54 seconds
===============================
Cheers,
Stan H.
===============================================
--------------------------------